OpenStudy (btaylor):

You have a 100 question multiple-choice test (multiple choice options A-D). Your statistics teacher says that every answer is the same (ie all the answers are A). Is there any way to get an expected value>25% on the test?

OpenStudy (btaylor):

If you guess that they're all A, then there is a 25% chance you will get a 100%. If you split up your answers, then you are guaranteed a 25%. Is there any combination with more than 25%?

OpenStudy (anonymous):

acully no , the only succes propertise u have is 25% by luck :P anless u are sure from one qs of the 100 !

OpenStudy (anonymous):

so it wud be like this \(\Huge\frac{1}{2}+\frac{1}{2}+...+\frac{1}{2}\text{ for 100 terms}\) \(\Huge \text{expected} = \frac{1}{2} ×100=50 % !!!\)

OpenStudy (anonymous):

unexpected to me to get that answer !

OpenStudy (anonymous):

Probably not ^.^ No matter how you split it... say, we let \[\Large x_n\] be the number of questions you answered with choice (n) (here being a, b, c, and d) etc Then the expected score would be \[\Large \mathbb{E}=\sum_{n\in \{a,b,c,d\}} x_n\left<p(x_n)\right>\] where \(\large p(x_n)\) is the probability that the correct answer is choice n. Anyway, that probability is presumably 25% for all four choices, so... \[\Large \mathbb E = \frac14x_a+\frac14x_b+\frac14x_c+\frac14x_d \] \[\Large \mathbb E = \frac14(x_a+x_b+x_c+x_d)\] But surely, the sum of all the \(\large x_n\) should be 100, since they should comprise all the questions... that's unless you intend to leave any questions blank... \[\Large \mathbb E = \frac14(100)=25\%\]

OpenStudy (btaylor):

thank you for that thorough reply!

OpenStudy (anonymous):

^.^ I do try

OpenStudy (anonymous):

@PeterPan \(\color{orange}{\text AWESOME ! !} \)